OpenStudy (anonymous):

how did 1/4 + 1/3 + 1/6 = 9/12 reduced to 3⁄4. 3⁄4 or 75%?? how did it get to 3/4? or 75%?

OpenStudy (anonymous):

9/12 Divide both top and bottom by the same simplest number, in this case is 3 If you divide boht sides by 3, 9/3 = 3 12/3 = 4 3/4 = 75%

OpenStudy (anonymous):

@dauspex you're the best! haha